OpenStudy (anonymous):

Solve sin2x = sqaureroot 2 cos x

OpenStudy (anonymous):

left side sin2x = 2sinx coax . square both side, you can have the solution

OpenStudy (zehanz):

Is it: \(\sin2x=\sqrt{2}\cos x\), or : \(\sin2x=\sqrt{2\cos x}\)?

OpenStudy (anonymous):

uhh sry its sin2x=\[\sqrt{2}\] cosx

OpenStudy (zehanz):

OK, so if you follow @Hoa's suggestion, you get: \[(2\sin x \cos x)^2=2 \cos^2x\]Can you do the next step?

OpenStudy (anonymous):

:) yes, i can hihi

OpenStudy (zehanz):

@Hoa: you an I can, but can @alexprip?

OpenStudy (anonymous):

Well I'm not exactly sure I have to find a way to remove the sin so I can have a quadratic equation right?

OpenStudy (zehanz):

YOu can split up into two different equations: one with cos²x and one with sin²x.

OpenStudy (zehanz):

\(4\sin^2x\cos^2x=2\cos^2x\) can be written as:\[4\sin^2x=2 \vee \cos^2=0\]

OpenStudy (zehanz):

That is because it is of this standard type: ab=ac <=> a=0 v b=c

OpenStudy (anonymous):

I'm confused..

OpenStudy (anonymous):

This one does it this way: http://answers.yahoo.com/question/index?qid=20090510225256AAJh74l

OpenStudy (zehanz):

standard type: ab=ac: there is a common factor on both sides. If a = 0 then you have 0b=0c, so 0 on both sides: This is ok, so a=0 is a solution. If a is NOT 0, then you can divide both sides by it, leaving b=c. That is why, as soon as you have an equation like ab=ac, you can replace it with a=0 v b=c.

OpenStudy (zehanz):

Now to use this on your equation: You have\[4\sin^2x \cos^2x=2\cos^2x \]This is \[\cos^2x \cdot 4\sin^2x = \cos^2 x \cdot 2\]Now you can see the ab=ac here: a = cos²x, b=4sin²x and c=2.

OpenStudy (anonymous):

I see but what do I then have to do next?

OpenStudy (zehanz):

OK, so now switch to a=0 v b=c:\[\cos^2x=0 \vee 4\sin^2x=2\]YOu can solve these two equations separately, they are easier than the whole thing!

OpenStudy (zehanz):

Do you know how to solve \(\cos^2x=0\)?

OpenStudy (anonymous):

Well I have to take a look at the unit circle right? \[\cos^{2}\] x = 0 => \[\frac{ \pi }{ 2 } and \frac{ 3\pi }{ 2 }\] right?

OpenStudy (anonymous):

Or am I completly wrong?

OpenStudy (zehanz):

You are right!

OpenStudy (zehanz):

Because cos²x=0 is the same as cosx=0.

OpenStudy (anonymous):

Great:) And for sin i divide by 4 and get sin squared = 1/2 ?

OpenStudy (zehanz):

Yes, so what should be the next step?

OpenStudy (anonymous):

sin squared = 1/2 \[\frac{ \pi }{ 6 } and \frac{ 5\pi }{ 6 }\] ?

OpenStudy (anonymous):

And that is about it?

OpenStudy (zehanz):

No, you have sin²x=1/2, so \[\sin x = \pm \sqrt{\frac{1}{2}}\pm \frac{1}{2}\sqrt{2}\]So you need to find x's that fit this.

OpenStudy (zehanz):

But it's your lucky day! These values \(\pm \frac{1}{2}\sqrt{2}\) belong to x's that are also easy to find in the unit circle...

OpenStudy (zehanz):

Here it is: